# Testable Flavored TeV-scale Resonant Leptogenesis with MeV-GeV Dark Matter in a Neutrinophilic 2HDM

### v1

@@ Abstract
We explore flavored resonant leptogenesis embedded in a neutrinophilic 2HDM. Successful leptogenesis is achieved by the very mildly degenerate two heavier right-handed neutrinos (RHNs) $N_{2}$ and $N_{3}$ with a level of only $\Delta M_{32}/M_{2}\sim\mathcal{O}(0.1\%-1\%)$. The lightest RHN, with a MeV–GeV mass, lies below the sphaleron freeze-out temperature and is stable, serving as a dark matter candidate. The model enables TeV-scale leptogenesis while avoiding the extreme mass degeneracy typically plagued conventional resonant leptogenesis. Baryon asymmetry, neutrino masses, and potentially dark matter relic density can be addressed within a unified, experimentally testable framework.
@@ I Introduction
The observed baryon asymmetry of the universe (BAU), $\eta_{B}\equiv\frac{n_{B}-n_{\bar{B}}}{n_{\gamma}}\bigg{|}_{0}\approx 6\times 10^{-10}$ [1, 2, 3], poses a significant challenge to the Standard Model (SM). While the SM satisfies the Sakharov conditions for baryogenesis [4, 5], it fails to generate sufficient asymmetry. Traditional thermal leptogenesis within the type-I seesaw model introduces heavy RHNs whose CP-violating, out-of-equilibrium decays generate a lepton asymmetry, which sphalerons partially convert to baryon asymmetry [6, 7]. However, the Davidson-Ibarra bound, valid for hierarchical RHNs, requires the lightest RHN to have a mass of $M_{1}\gtrsim 10^{9}$ GeV for sufficient asymmetry to be generated [8, 9, 7]. This heavy scale exacerbates the hierarchy problem of the SM [10] and introduces cosmological challenges, such as the gravitino problem in supersymmetric extensions [11, 12]. Furthermore, this scale is far beyond the reach of any foreseeable experiments, making experimental verification difficult, if not impossible.
Resonant leptogenesis alleviates these issues by enhancing CP asymmetry through the interference between nearly degenerate RHN states [13, 14, 15, 16, 17]. In this scenario, the self-energy contribution to CP asymmetry becomes resonantly enhanced when the mass splitting between two RHNs is comparable to their decay widths, $\Delta M\sim\Gamma$. This allows successful leptogenesis at TeV scales. However, achieving such near-degeneracy requires extreme fine-tuning of the RHN masses, $\Delta M/M\sim 10^{-9}$, which is viewed as unnatural.
The $\nu$2HDM (neutrinophilic Two-Higgs Doublet Model) offers an alternative by extending the SM+RHNs with a second Higgs doublet having a small vacuum expectation value (VEV) and coupling exclusively to neutrinos. The smaller VEV reduces the equilibrium neutrino masses, increasing the CP asymmetry and allowing leptogenesis to occur at lower scales [18, 19, 20, 21]. However, $\nu$2HDM often suffer from strong $\Delta L=2$ washout processes, which impose a lower limit of $M_{N}\gtrsim 10^{5}$ GeV, making experimental verification still challenging.
@@
To address these challenges, this paper proposes a model combining the approaches of resonant leptogenesis and $\nu$2HDM. Inspired by neutrino mass pattern model building, we also introduce a RHN mass hierarchy: $M_{1}\ll T_{\text{spha}}<M_{2}\sim M_{3}$. Upon accounting for flavor effects, $N_{2}$ and $N_{3}$ enable TeV-scale resonant leptogenesis with only mild mass degeneracy breaking, $\Delta M_{32}/M_{2}\sim\mathcal{O}(0.1\%-1\%)$. If this mass pattern arises from a fundamental symmetry breaking, the breaking of degeneracy directly relates to the scale of $N_{1}$ [22], implying $M_{1}\sim\mathcal{O}(\text{MeV-GeV})\ll T_{\text{spha}}$ in our scenario. Since sphalerons remain in equilibrium until $T=131.7$ GeV [23], the lightest $N_{1}$ decouples from thermal leptogenesis and is stable. Thus, the model naturally addresses both BAU and the DM relic density, a feature absent in most thermal leptogenesis models. Furthermore, all model parameters lie in experimentally testable range, ensuring verification in the near future.
@@ II Neutrino Masses
We extend the SM by a Type-I 2HDM, where the second Higgs doublet, $\Phi_{2}$, couples exclusively to neutrinos. Both $\Phi_{1}$ and $\Phi_{2}$ have hypercharge $+1$. A discrete $Z_{2}$ symmetry is imposed to suppress flavor-changing neutral currents at tree level, with $\Phi_{1}$ and SM fermions being $Z_{2}$-even, while $\Phi_{2}$ and the RHNs are $Z_{2}$-odd.
In this framework, $\Phi_{1}$ generates masses for SM fermions, while $\Phi_{2}$, with a small VEV $v_{2}$, facilitates the seesaw mechanism for neutrino masses. The SM VEV is $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\approx 246$ GeV, and we consider $v_{2}\ll v_{1}$ such that $\tan\beta=v_{1}/v_{2}\gg 1$. In other types of 2HDM, such as Type II, X, and Y, the Yukawa couplings hit a Landau pole for large $\tan\beta$, limiting the parameter space [24]. The Type-I 2HDM avoids this issue and remains perturbative up to the Planck scale. The Yukawa interactions are described by
@@
where $U_{\text{PMNS}}$ is the PMNS matrix, and $R$ is a complex orthogonal matrix parameterized by three complex rotation angles (or six real angles) $z_{i}=x_{i}+iy_{i}$ ($i=1,2,3$).
@@
This mass pattern is crucial: $N_{1}$ does not participate in thermal leptogenesis because when $N_{1}$ becomes thermally important ($T\sim M_{1}$), the temperature of the universe is already lower than $T_{\text{spha}}$, at which point all baryon asymmetry converted from lepton asymmetry is frozen. Additionally, in the decoupling limit of 2HDM, where the observed SM-like Higgs boson corresponds to the light neutral CP-even state [32], $m_{H},m_{A},m_{H^{\pm}}\gg m_{h}\sim 125$ GeV $\gg M_{1}$, there are no decay channels for $N_{1}$. Thus, $N_{1}$ is stable and serves as a dark matter candidate. Out model thus address both the BAU (through the mildly degenerate $N_{2}$ and $N_{3}$) and the observed DM relic density (through the stable $N_{1}$) without the need to introduce additional fields or new mechanisms.
@@ IV Numerical Analysis
The results are displayed as color contour maps to reveal patterns and correlations among the scanned parameters, particularly the mixing angles and phases. For instance, Fig. 2 shows the contour on the plane of $\alpha_{21}$ versus $\delta_{\text{CP}}$ under the IO spectrum. The color represents the magnitude of the final baryon asymmetry $|\eta_{B}|$. The observed $\eta_{B}$ corresponds to the boundary of red and orange. A clear preference for $\alpha_{21}\in[45^{\circ},315^{\circ}]\cup[405^{\circ},675^{\circ}]$ is evident while $\delta_{\text{CP}}$ shows no strong preference. $\alpha_{31}$ was also found to be irrelevant to the baryon asymmetry. A complete set of such scan contours is discussed and presented in Appendix A. For NO, our results indicate that both Dirac and Majorana phases have negligible impact.
Our scans also show that $M_{1}$ has no influence on the baryon asymmetry. However, the lightest light neutrino mass $m_{1(3)}$ exhibits a threshold for generating sufficient baryon asymmetry. For $\tan\beta=225$, $m_{1(3)}\lesssim 10^{-3}$ eV is required. This behavior arises because a smaller $v_{2}$ reduces the equilibrium neutrino mass $m_{*}$, necessitating a reduction in the effective neutrino mass $\tilde{m}_{2\alpha}$ to counteract enhanced washout. Since $\tilde{m}_{2\alpha}$ does not depend on $v_{2}$, the suppression of the lightest neutrino mass becomes essential to partially balance the washout factor $K$ in Eq. (9) for fixed $M_{2}$.
@@
Figure 2: BAU dependence as a contour on $\delta_{\text{CP}}$ and $\alpha_{21}$ plane for IO neutrino spectrum. The color corresponds to the magnitude of $\eta_{B}$ survived. The boundary of red and orange corresponds to the observed value.
The contours in appendix reveal highly structured patterns and correlations among the CI mixing angles for successful BAU. Notably, these patterns tend to select specific flavors responsible for leptogenesis, depending on the neutrino mass ordering. For NO, successful leptogenesis has a strong taste of the $e$ flavor for generating the excess, while for IO, it has a distaste of the $e$ flavor and instead nearly equally favors $\mu$ or $\tau$ flavors. These distinct patterns may suggest a potential underlying flavor symmetry that constrains the mixing angles. We leave the possibility that the same symmetry could also govern the RHN mass patterns. Further investigation into this flavor symmetry could provide valuable insights into the interplay between mixing angles and mass patterns in the RHN sector. The contours also suggest IO case allows a richer parameter space than the NO case.
Based on the scan results, we identified targeted benchmark points (BMs) to demonstrate the viability of leptogenesis within our model. These points satisfy neutrino oscillation experimental constraints, generate sufficient BAU, and lie outside the strong $\Delta L=2$ washout region shown in Fig. 1. The BMs were chosen to reflect distinct levels of $M_{2}$, mass degeneracy levels, and $\tan\beta$: either a) small $M_{2}$ with high degeneracy (still mild in general terms), b) large $M_{2}$ with low degeneracy, or c) a trade-off between these two directions. The BMs are summarized in Table 1, while a complete set of parameters for these BMs is given in Appendix B.
@@ V Conclusion
In this study, we have demonstrated that TeV-scale leptogenesis can be achieved with only a very mild mass degeneracy between the RHNs $N_{2}$ and $N_{3}$. This degeneracy can naturally arise, for instance, from a flavor symmetry whose soft breaking results in the required mass pattern. Such a flavor symmetry might also explain the structured patterns observed in the scan of Casas-Ibarra rotation angles, as shown in Appendix A. A UV-complete realization of this scenario is left for future investigation. Importantly, in such a scenario, the mass of the lightest right-handed neutrino, $N_{1}$, is naturally below both the sphaleron freeze-out temperature $T_{\text{spha}}$ and the SM-like Higgs boson mass $m_{h}$.
Traditional leptogenesis models often face challenges in simultaneously addressing the BAU and the dark matter relic density, as they typically rely on the decay of the lightest RHN, $N_{1}$, or risk $N_{1}$ washing out the asymmetry generated by $N_{2}$ and $N_{3}$. Our model circumvents this issue by placing $N_{1}$ below $T_{\text{spha}}$, rendering it inactive during leptogenesis. Furthermore, since $N_{1}$ is lighter than the SM-like Higgs boson, it lacks tree-level decay channels, making it a viable DM candidate. Thus, both BAU and the DM relic density can be accounted for.
@@ Acknowledgements.
Appendix A Scan Contours of Parameters
(a) $\tan\beta=225$
(b) $\tan\beta=100$
[…]
(a) $\tan\beta=225$
(b) $\tan\beta=100$
[…]
This appendix presents the detailed scan contours for the Casas-Ibarra mixing angles $x_{i}$ and $y_{i}$, the lightest heavy and light neutrino masses, the Dirac phase $\delta_{\text{CP}}$, and the Majorana phase $\alpha_{21}$ for both NO and IO neutrino spectra. These plots reveal structured correlation patterns among the parameters that lead to successful baryon asymmetry generation.
Fig. 4 shows the scan results for the NO spectrum. Panel (a) depicts the contours for $\tan\beta=225$, while panel (b) shows the results for $\tan\beta=100$ to illustrate the impact of varying $\tan\beta$. As noted in the main text, varying $\tan\beta$ only shifts the boundaries of the amount of BAU on the contour plots without altering the internal correlations and patterns among the mixing angles and phases. In general for NO,
  • $|y_{2}|$ prefers larger values.
  • $|x_{3}|$ clusters near $90^{\circ}$,
  • $|y_{3}|$ prefers small values.
  • $x_{2}$, $\delta_{\text{CP}}$, $\alpha_{21}$, and $\alpha_{31}$ exhibit no preference for leptogenesis.
Figure 5 presents the scan results for the IO spectrum. We observed that
  • $|x_{1}|$ and $|y_{1}|$ tend to cluster near $0^{\circ}$.
  • $|x_{2}|$ avoids $90^{\circ}$.
  • Unlike NO, $\alpha_{21}$ shows a clear preference for $[45^{\circ},315^{\circ}]\cup[405^{\circ},675^{\circ}]$, while $x_{3}$, $\delta_{\text{CP}}$ and $\alpha_{31}$ remain irrelevant for leptogenesis.
Appendix B Parameters for Leptogenesis Benchmark Points

### v2

@@ Abstract
We explore flavored resonant leptogenesis embedded in a neutrinophilic 2HDM. Successful leptogenesis is achieved by the very mildly degenerate two heavier right-handed neutrinos (RHNs) $N_{2}$ and $N_{3}$ with a level of only $\Delta M_{32}/M_{2}\sim\mathcal{O}(0.1\%-1\%)$. The lightest RHN, with a MeV–GeV mass, lies below the sphaleron freeze-out temperature and is stable, serving as a dark matter candidate. The model enables TeV-scale leptogenesis while avoiding the extreme mass degeneracy typically plagued conventional resonant leptogenesis. Baryon asymmetry, neutrino masses, and potentially even dark matter relic density can be addressed within a unified, experimentally testable framework.
@@ I Introduction
The observed baryon asymmetry of the universe (BAU), $\eta_{B}\equiv\frac{n_{B}-n_{\bar{B}}}{n_{\gamma}}\bigg{|}_{0}\approx 6\times 10^{-10}$ [1, 2, 3], poses a significant challenge to the Standard Model (SM). While the SM satisfies the Sakharov conditions for baryogenesis [4, 5], it fails to generate sufficient asymmetry. In particular, more baryon number violation, more CP violation, and a departure from thermal equilibrium unlike the one in the SM are needed. Traditional thermal leptogenesis within the type-I seesaw model introduces heavy RHNs whose CP-violating, out-of-equilibrium decays generate a lepton asymmetry, which sphalerons partially convert to baryon asymmetry [6, 7]. However, the Davidson-Ibarra bound, valid for hierarchical RHNs, requires the lightest RHN to have a mass of $M_{1}\gtrsim 10^{9}$ GeV for sufficient asymmetry to be generated [8, 9, 7]. This heavy scale exacerbates the hierarchy problem of the SM [10] and introduces cosmological challenges, such as the gravitino problem in supersymmetric extensions [11, 12]. Furthermore, this scale is far beyond the reach of any foreseeable experiments, making experimental verification difficult, if not impossible.
Resonant leptogenesis alleviates these issues by enhancing CP asymmetry through the interference between nearly degenerate RHN states [13, 14, 15, 16, 17, 18, 19]. In this scenario, the self-energy contribution to CP asymmetry becomes resonantly enhanced when the mass splitting between two RHNs is comparable to their decay widths, $\Delta M\sim\Gamma$. This allows successful leptogenesis at TeV scales. However, achieving such near-degeneracy requires extreme fine-tuning of the RHN masses, $\Delta M/M\sim 10^{-9}$, which is viewed as unnatural.
The neutrinophilic Two-Higgs Doublet Model ($\nu$2HDM) offers an alternative by extending the SM+RHNs with a second Higgs doublet having a small vacuum expectation value (VEV) and coupling exclusively to neutrinos. The smaller VEV reduces the equilibrium neutrino masses, increasing the CP asymmetry and allowing leptogenesis to occur at lower scales [20, 21, 22, 23]. However, $\nu$2HDM often suffers from strong $\Delta L=2$ washout processes. Together with the Davidson-Ibarra bound, which still applies to these hierarchical leptogenesis models with a modified VEV, they impose a lower limit on the RHN mass of $M_{N}\gtrsim 10^{5}$eV [21], making experimental verification still challenging.
@@
To address these challenges, this paper proposes a model combining the approaches of resonant leptogenesis and $\nu$2HDM. Inspired by neutrino mass pattern model building, we also introduce a RHN mass hierarchy: $M_{1}\ll T_{\text{spha}}<M_{2}\sim M_{3}$. Upon accounting for flavor effects, $N_{2}$ and $N_{3}$ enable TeV-scale resonant leptogenesis with only mild mass degeneracy breaking, $\Delta M_{32}/M_{2}\sim\mathcal{O}(0.1\%-1\%)$. If this mass pattern arises from a fundamental symmetry breaking, the breaking of degeneracy directly relates to the scale of $N_{1}$ [24], implying $M_{1}\sim\mathcal{O}(\text{MeV-GeV})\ll T_{\text{spha}}$ in our scenario. Since sphalerons remain in equilibrium until $T=131.7$ GeV [25], the lightest $N_{1}$ decouples from thermal leptogenesis and is stable. Thus, the model potentially addresses both BAU and the DM relic density, a feature absent in most thermal leptogenesis models. Furthermore, all model parameters lie in experimentally testable range, ensuring verification in the near future.
@@ II Neutrino Masses
We extend the SM by a second Higgs doublet. The Higgs sector couples to SM fermions in a Type-I 2HDM-like manner. We also introduce RHNs, which couple exclusively to the second Higgs doublet $\Phi_{2}$. Both $\Phi_{1}$ and $\Phi_{2}$ have hypercharge $+1$. A discrete $Z_{2}$ symmetry is imposed to suppress flavor-changing neutral currents at tree level, with $\Phi_{1}$ and SM fermions being $Z_{2}$-even, while $\Phi_{2}$ and the RHNs are $Z_{2}$-odd. Such scheme is often dubbed as $\nu$2HDM [20, 21, 22, 23].
In this framework, $\Phi_{1}$ generates masses for SM fermions, while $\Phi_{2}$, with a small VEV $v_{2}$, facilitates the seesaw mechanism for neutrino masses. The SM VEV is $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\approx 246$ GeV, and we consider $v_{2}\ll v_{1}$ such that $\tan\beta=v_{1}/v_{2}\gg 1$. If one starts from other types of 2HDM, such as Type II, X, and Y, the Yukawa couplings hit a Landau pole for large $\tan\beta$, limiting the parameter space [26]. The $\nu$2HDM avoids this issue and remains perturbative up to the Planck scale. The Yukawa interactions are described by
@@
where $U_{\text{PMNS}}$ is the standard PMNS matrix with Majorana phases $\alpha_{21}$ and $\alpha_{31}$:
$U_{\text{PMNS}}=\hat{U}\cdot\text{diag}(1,e^{i\frac{\alpha_{21}}{2}},e^{i\frac{\alpha_{31}}{2}}),$ (4)
and $R$ is a complex orthogonal matrix parameterized by three complex rotation angles (or six real angles) $z_{i}=x_{i}+iy_{i}$ ($i=1,2,3$).
@@
This mass pattern is crucial: $N_{1}$ does not participate in thermal leptogenesis because when $N_{1}$ becomes thermally important ($T\sim M_{1}$), the temperature of the universe is already lower than $T_{\text{spha}}$, at which point all baryon asymmetry converted from lepton asymmetry is frozen. Additionally, in the decoupling limit of 2HDM, where the observed SM-like Higgs boson corresponds to the light neutral CP-even state [34], $m_{H},m_{A},m_{H^{\pm}}\gg m_{h}\sim 125$ GeV $\gg M_{1}$, there are no two-body tree-level decay channels for $N_{1}$. Three-body decays are, in principle, possible but are very suppressed via an off-shell heavy Higgs or the SM-like neutral Higgs. For example, if mediated by a heavy charged Higgs (e.g. $N_{1}\to\nu\,\ell^{+}\ell^{-}$), the decay is highly suppressed by the off-shell propagator, and the decay width can be approximated as $\Gamma_{N_{1}}\sim\frac{|y_{\nu}|^{4}}{192\pi^{3}}\frac{M_{1}^{5}}{m_{H}^{4}}$. For TeV-scale heavy Higgs bosons and a sub-GeV $N_{1}$, this decay width is much smaller than the Hubble constant, ensuring that $N_{1}$ is effectively stable and contribute to the relic density. Similarly, three-body decays mediated by the SM-like neutral Higgs (e.g., $N_{1}\to 3\nu$) are suppressed by extremely tiny couplings between the SM-like Higgs and the SM neutrinos. While the decays are highly suppressed, the production of $N_{1}$ is very efficient through 2$\to$2 scattering processes when the mediators are on-shell, ensuring that a sufficient relic abundance is produced. Thus, $N_{1}$ can serve as a dark matter candidate. Our model, therefore, has the potential to simultaneously address the BAU (via the mildly degenerate $N_{2}$ and $N_{3}$) and the observed dark matter relic density (through the stable $N_{1}$) without introducing additional fields or mechanisms.
@@ IV Numerical Analysis
The results are displayed as color contour maps to reveal patterns and correlations among the scanned parameters, particularly the mixing angles and phases. For instance, Fig. 2 shows the contour on the plane of $\alpha_{21}$ versus $\delta_{\text{CP}}$ under the NO and IO spectrum, respectively. The color represents the magnitude of the final baryon asymmetry $|\eta_{B}|$. The observed $\eta_{B}$ corresponds to the boundary of red and orange. For the NO spectrum, the generated BAU does not depend on either Dirac or Majorana CP phases. For the IO spectrum, a clear preference for $\alpha_{21}\in[45^{\circ},315^{\circ}]\cup[405^{\circ},675^{\circ}]$ is evident while $\delta_{\text{CP}}$ shows no strong preference. In both spectrum, $\alpha_{31}$ was also found to be irrelevant to the baryon asymmetry.
(a) NO
(b) IO
Figure 2: BAU dependence as a contour on $\delta_{\text{CP}}$ and $\alpha_{21}$ plane. The color corresponds to the magnitude of $\eta_{B}$ survived. The boundary of red and orange corresponds to the observed value.
(a) NO
(b) IO
Figure 3: BAU dependence as a contour on the mass of the lightest LH neutrino $m_{1(3)}$ and the mass of the lightest RH neutrino $M_{1}$.
Fig. 3 shows the contour scan of the lightest left-handed neutrino mass $m_{1(3)}$ and right-handed neutrino mass $M_{1}$. Our scans show that $M_{1}$ has no influence on the baryon asymmetry. However, the lightest light neutrino mass $m_{1(3)}$ exhibits a threshold for generating sufficient baryon asymmetry. For $\tan\beta=225$, $m_{1(3)}\lesssim 10^{-3}$ eV is required. This behavior arises because a smaller $v_{2}$ reduces the equilibrium neutrino mass $m_{*}$, necessitating a reduction in the effective neutrino mass $\tilde{m}_{2\alpha}$ to counteract enhanced washout. Since $\tilde{m}_{2\alpha}$ does not depend on $v_{2}$, the suppression of the lightest neutrino mass becomes essential to partially balance the washout factor $K$ in Eq. (10) for fixed $M_{2}$.
(a) $\tan\beta=225$
(b) $\tan\beta=100$
[…]
(a) $\tan\beta=225$
(b) $\tan\beta=100$
[…]
@@
A detailed scan contours for the Casas-Ibarra mixing angles $x_{i}$ and $y_{i}$ for both NO and IO neutrino spectra is also performed. Fig. 4 shows the scan results for the NO spectrum. Panel (a) depicts the contours for $\tan\beta=225$, while panel (b) shows the results for $\tan\beta=100$ to illustrate the impact of varying $\tan\beta$. As mentioned above, varying $\tan\beta$ only shifts the boundaries of the amount of BAU on the contour plots without altering the internal correlations and patterns among the mixing angles and phases. In general, for NO,
  • $|y_{2}|$ prefers larger values.
  • $|x_{3}|$ clusters near $90^{\circ}$,
  • $|y_{3}|$ prefers small values.
  • $x_{2}$, $\delta_{\text{CP}}$, $\alpha_{21}$, and $\alpha_{31}$ do not exhibit a preference for leptogenesis.
Fig. 5 presents the scan results for the IO spectrum. We observed that
  • $|x_{1}|$ and $|y_{1}|$ tend to cluster near $0^{\circ}$.
  • $|x_{2}|$ avoids $90^{\circ}$.
  • Unlike NO, $\alpha_{21}$ shows a clear preference for $[45^{\circ},315^{\circ}]\cup[405^{\circ},675^{\circ}]$, while $x_{3}$, $\delta_{\text{CP}}$, and $\alpha_{31}$ remain irrelevant for leptogenesis.
These contours reveal highly structured patterns and correlations among the CI mixing angles for successful BAU. Notably, these patterns actually tend to select specific flavors responsible for leptogenesis, depending on the neutrino mass ordering. For NO, successful leptogenesis has a strong taste of the $e$ flavor for generating the excess, while for IO, it has a distaste of the $e$ flavor and instead nearly equally favors $\mu$ or $\tau$ flavors. These distinct patterns may suggest a potential underlying flavor symmetry that constrains the mixing angles. We leave the possibility that the same symmetry could also govern the RHN mass patterns. Further investigation into this flavor symmetry could provide valuable insights into the interplay between mixing angles and mass patterns in the RHN sector. The contours also suggest the IO case allows a richer parameter space than the NO case.
Based on the scan results, we identified targeted benchmark points (BMs) to demonstrate the viability of leptogenesis within our model. These points satisfy neutrino oscillation experimental constraints, generate sufficient BAU, and lie outside the strong $\Delta L=2$ washout region shown in Fig. 1. The BMs were chosen to reflect distinct levels of $M_{2}$, mass degeneracy levels, and $\tan\beta$: either a) small $M_{2}$ with high degeneracy (still mild in general terms), b) large $M_{2}$ with low degeneracy, or c) a trade-off between these two directions. The BMs are summarized in Table 1, while a complete set of parameters for these BMs can be found in the appendix A.
@@ V Conclusion
In this study, we have demonstrated that TeV-scale leptogenesis can be achieved with only a very mild mass degeneracy between the RHNs $N_{2}$ and $N_{3}$. This degeneracy can naturally arise, for instance, from a flavor symmetry whose soft breaking results in the required mass pattern. Such a flavor symmetry might also explain the structured patterns observed in the scan of Casas-Ibarra rotation angles. A UV-complete realization of this scenario is left for future investigation. Importantly, in such a scenario, the mass of the lightest right-handed neutrino, $N_{1}$, is naturally below both the sphaleron freeze-out temperature $T_{\text{spha}}$ and the SM-like Higgs boson mass $m_{h}$.
Traditional leptogenesis models often face challenges in simultaneously addressing the BAU and the dark matter relic density, as they typically rely on the decay of the lightest RHN, $N_{1}$, or risk $N_{1}$ washing out the asymmetry generated by $N_{2}$ and $N_{3}$. Our model circumvents this issue by placing $N_{1}$ below $T_{\text{spha}}$, rendering it inactive during leptogenesis. Furthermore, since $N_{1}$ is lighter than the SM-like Higgs boson, it lacks 2-body tree-level decay channels while 3-body tree-level is very suppressed, making it a viable DM candidate. Thus, both BAU and potentially the DM relic density can be accounted for.
@@ Acknowledgements.
Appendix A Parameters for Leptogenesis Benchmark Points